OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

part 1 means your equation is balanced.

OpenStudy (foolaroundmath):

Part 1: It means that the system of equations is always true regardless of what values the unknown variables take. Thus, you get infinite number of solutions for that system Part 2: Add the two equaions: 3y - 2x + y+ 2x = 11 + 9 => 4y = 20 => y = 5 Substituting in second equation, 5 + 2x = 9 => 2x = -4 => x = -2 So, (x,y) = (-2,5)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Part 2: Add the two equaions: 3y - 2x + y+ 2x = 11 + 9 => 4y = 20 => y = 5 Substituting in second equation, 5 + 2x = 9 => 2x = 4 => x = 2 So, (x,y) = (2,5)

OpenStudy (anonymous):

Part 2: Add the two equaions: 3y - 2x + y+ 2x = 11 + 9 => 4y = 20 => y = 5 Substituting in second equation, 5 + 2x = 9 => 2x = 4 => x = 2 So, (x,y) = (2,5)